OpenStudy (anonymous):

If the parallel sides of a trapezoid contained by the lines y = -x + 5 and y = -x - 9, what equation represents the line contained by the midsegment?

OpenStudy (anonymous):

A. y = -x – 2 B. y = x + 2 D. y = -x + 7 I know it isn't C